OpenStudy (anonymous):

which statement is true about the root of the equation below? square root of x^2-5x+5=1

OpenStudy (anonymous):

1)The only root is 1 2) The only root is 4 3)Both 1 and 4 are roots 4)Neither 1 nor 4 is a root

OpenStudy (solomonzelman):

I would solve it. \[\huge\color{red}{ x^2-5x+5=1 } \]\[\huge\color{red}{ x^2-5x+5 } +1.25=\huge\color{red}{ 1 } +1.25\]\[\huge\color{red}{ x^2-5x+6.25=2.25 } \] and we know that \[1.5^2=2.25~~~~~~~~and~~~~~~~~~2.5^2=6.25\]can tyou solve it from here?

OpenStudy (anonymous):

i got 1 and 4 for the roots

OpenStudy (solomonzelman):

YUP!

OpenStudy (solomonzelman):

Good work.

OpenStudy (anonymous):

thanks!!

OpenStudy (solomonzelman):

Anytime, the reason I thought of completing the square because it's easier to do it without a paper.

OpenStudy (solomonzelman):

I made you yellow just now, CONGRATS on your new color!

OpenStudy (anonymous):

Aww.. thank youuu :D

OpenStudy (kc_kennylau):

Wow congratulations :D

OpenStudy (kc_kennylau):

@SolomonZelman but factorizing x^2-5x-4=0 is easier?

OpenStudy (kc_kennylau):

+4*

OpenStudy (solomonzelman):

Idk, I like completing the square, also it's good to remember that, \[\huge\color{blue}{ 12^2=144 } ~~~~\huge\color{green}{ 13^2=169 } ~~~~\huge\color{red}{ 14^2=196 } \]\[\huge\color{goldenrod}{ 15^2=225 } ~~~....~~~\huge\color{magenta }{ 25^2=625 } \]

OpenStudy (solomonzelman):

You do have the point though.

OpenStudy (kc_kennylau):

I see. So different people have different "easiet" approaches :)